OpenStudy (anonymous):

The ratio of the heights of two similar pyramids is 4:7. The volume of the smaller pyramid is 1,331cm^3. To the nearest whole number ,what is the volume of the larger pyramid? A) 154cm^3 B) 7,133cm^3 C) 14,641cm^3 D) 18,643cm^3

OpenStudy (dumbcow):

ratio of volumes is 4^3/7^3 = 64/343 64/343 = 1331/X cross multiply and solve for X

OpenStudy (anonymous):

so is it B?